# Extremely luminous optical afterglow of a distant and energetic gamma-ray burst GRB 230204B

### v1

@@ 2.4.2 BOOTES robotic telescope
BOOTES followed up GRB 230204B with three of the six 60 cm robotic telescopes, namely the ones at the BOOTES-4/MET station in Lijiang Astronomical Observatory (Yunnan, China), at BOOTES-5/JGT station in Observatorio Astronómico Nacional of San Pedro Mártir (Baja California, México) and at the BOOTES-2/TELMA station in La Mayora (Málaga, Spain). Following the MAXI trigger, the BOOTES-4/MET telescope performed two epoch observations at 2023-02-04 22:19:12 UT and 2023-02-05 17:42:04 UT. The BOOTES-2/TELMA telescope and the BOOTES-5/JGT telescope also executed follow-up observations at 2023-02-05 04:16:55 UT and 2023-02-05 07:43:45 UT, respectively. A series of images were obtained using the clear filter with exposures of 1 sec, 5 sec, and 10 sec at BOOTES-4/MET telescope during the first epoch and the optical afterglow was clearly detected during the decaying phase. However, during the following multi-filters (Sloan- and clear filters, 60 sec exposure) monitoring, the optical afterglow was not detected on the stacked images for each epoch (as they were not as deep as the clear filter exposures). All the images were processed with bias, dark subtraction, and flat-field correction using custom IRAF routines (Gupta et al., 2021b). Photometry was carried out using the standard IRAF package and the images were calibrated with nearby comparison stars from the USNO-B1.0 and SDSS catalogs.
@@ 2.4.3 DOT
We initiated observations using the DOT (Pandey, 2016; Gupta et al., 2024a) on 2023-02-06 at 01:49:47.939 UT, approximately 1.16 days after the trigger. Multiple frames with an exposure time of 180 sec each were observed in the , , and bands, and continuous observations were conducted over four consecutive nights using the same filters. For image preprocessing, we employed the standard IRAF framework, and DAOPHOT II was used to determine magnitudes from the cleaned files. We stacked the images for each night in the same filter to increase the signal-to-noise ratio. PSF photometry was performed on the final stacked images (Gupta et al., 2022b, c). Due to the high moon phase during the burst observation, we only detected GRB 230204B in the and -band images from the first night. For the remaining filters, we could only establish upper limits. Photometric calibration is performed using the standard stars from the Pan-STARRS catalog.
@@ 2.4.4 Swift/UVOT
The Swift/UVOT began settled observations of the field of GRB 230204B 80.6 ks after the MAXI trigger (Serino et al., 2023). No optical afterglow was detected by UVOT. Photometry was obtained from co-added UVOT image mode data. The source counts were extracted from the UVOT stacked images using a source region of 5 arcsec radius. Background counts were extracted using an aperture of 20 arcsec in radius located in a source-free region near the GRB. We use the Swift tool uvotsource to obtain the background subtracted count rates using source and background apertures of 5 and 20 arcseconds in radius, respectively. They were converted to magnitude using the UVOT photometric zero points (Breeveld et al., 2011; Poole et al., 2008).
@@ 2.4.5 Radio Emission
We observed GRB 230204B on 8 February 2023 (3.6 days post-burst), with the Australia Telescope Compact Array (ATCA, Project code CX527) at a wide range of frequencies centered on 5.5, 9.0, 16.7, 21.2, 33 and 35 GHz, each with a 2048 MHz-wide band. We reduced the visibility data using standard routines in MIRIAD (Sault et al., 1995). We used a combination of manual and automatic RFI flagging before calibration, conducted with MIRIAD tasks uvflag and pgflag, respectively. We used PKS 193463 to determine the bandpass response for frequency bands 5.5 and 9 GHz, and PKS 0727115 as the bandpass calibrator for 16.7, 21.2, 33 and 35 GHz bands. We used PKS 193463 to calibrate the flux density scale for all frequency bands. We used PKS 1256220 to calibrate the time-variable complex gains for all epochs and frequency bands. After calibration, we inverted the visibilities using a robust weighting of 0.5 and then used the CLEAN algorithm (Clark, 1980) on the target source field using standard MIRIAD tasks INVERT, CLEAN and RESTOR to obtain the final images.
We detect the radio counterpart at 16.7 and 21.2 GHz at a position consistent with the BOOTES/MASTER/GIT optical counterpart position (Swain et al., 2023). For all other frequencies, we report 5-sigma upper limits (see Table 7 of appendix). The GRB position was also observed with the Australian SKA Pathfinder by the Rapid ASKAP Continuum Survey (RACS; McConnell et al., 2020) on 2023 July 1 and 2024 Jan 5. We find no radio detections in preliminary data and report 5-sigma upper limits of 1.5 mJy. There were also no archival radio detection pre-burst within 1 arcminute in surveys: the National Radio Astronomy Observatory VLA Sky Survey (Condon et al., 1998), the Sydney University Molonglo Sky Survey (Mauch et al., 2003), the Very Large Array Sky Survey (Lacy et al., 2020) or the RACS survey. We measured a VLASS 5-sigma upper limit of 0.69 mJy at 3 GHz.
@@ 2.4.6 Swift BAT Survey Data
Swift/XRT did not observe the location of X-ray afterglow of GRB 230204B until 91.5 ks post-burst. Therefore to search for the hard X-ray early afterglow, we used the BatAnalysis python package (Parsotan et al., 2023) to analyze Swift BAT survey data from 2023-02-03 to 2023-02-11. We analyzed survey data, where the location of the GRB had at least a partial coding fraction of on the BAT detector plane. As expected, the GRB afterglow was not detected in the BAT survey data, which is typical since hard X-ray afterglows have never been detected at such late times. However, we were able to place 5 upper limits on the emission in the 14-195 keV energy band (shown in Figure 2). We additionally conducted a mosaiced analysis of the BAT survey data in 1-day time bins, for two days after the trigger time, and were able to place 5 flux upper limits on the emission in the 14-195 keV energy band where the flux of the GRB afterglow is erg/sec/cm2.
@@ 3 Results
3.1 Prompt emission temporal and spectral characterize
For the analysis of Fermi GBM data, we utilized the Fermi GBM Data Tools222https://fermi.gsfc.nasa.gov/ssc/data/analysis/gbm/gbm_data_tools/gdt-docs/ to extract temporal information. We selected specific NaI (NaI 7 and NaI 8) and BGO (BGO 1) detectors for the analysis based on their optimal viewing angles relative to the GRB’s position to maximize the signal-to-noise ratio. For the NaI detectors, we focused on two energy bands: 8-30 keV and 50-300 keV, allowing us to explore both low and high-energy photon populations. For BGO 1, two bands were used: 0.3-1 MeV and 1-40 MeV, to investigate higher-energy photons. The prompt emission background-subtracted light curve of GRB 230204B, obtained using Fermi GBM across different energy channels, reveals a complex multi-pulse structure, as shown in Figure 3. This GRB exhibits four distinct emission episodes, each separated by periods of quiescence, with the shaded regions highlighting the intervals selected for time-integrated spectral analysis. The energy-dependent light curves show that the temporal structure and intensity vary significantly across different energy bands, reinforcing the diverse nature of GRB light curves. Notably, the hardness ratio (HR) demonstrates evolution over time, indicating spectral changes during the burst. The inset plot shows the MAXI light curve, covering the 2-20 keV energy range, further subdivided into 2-4 keV, 4-10 keV, and 10-20 keV. Although MAXI detected only part of the GRB, its observations provide valuable insights into the lower-energy emission and suggest a very softer nature of the last pulse of GRB 230204B.
We analyzed the time-integrated (T0-0.80 to T0 +229.69 sec) Fermi spectrum of GRB 230204B using the Multi-Mission Maximum Likelihood framework (3ML) to evaluate its spectral characteristics. We explore several phenomenological models, including the Band function, Cutoff power-law (CPL), and combinations with a Blackbody component, which is theoretically expected to originate from the photospheric emission of GRBs. For each model, we employed Bayesian spectral fitting to obtain the Deviance Information Criterion (DIC) values, allowing us to systematically compare the model fits. The results indicate that the Band function provides a statistically better fit to the time-integrated spectrum compared to the CPL alone, as indicated by the DIC values. Furthermore, the addition of a Blackbody component alongside the Band function yielded an even lower DIC value, suggesting that this model combination better describes the observed spectral features.
@@ 3.1 Prompt emission temporal and spectral characterize
GRBs are known for their significant spectral evolution and high variability, with emission properties often varying notably across individual pulses within multi-pulse bursts. Li et al. (2021) conduct a comprehensive time-resolved Bayesian spectral analysis of well-defined GRB pulses observed by Fermi. They observe a trend toward spectral softening over time, largely reflected in the gradual decrease of the low-energy power-law index. We compared the time-resolved spectral analysis of GRB 230204B with the sample studied by Li et al. (2021). We also observe that the low-energy power-law indices of GRB 230204B become softer over time (hard to soft), with a distinct spectrum evolution from the first to later episodes (see Figure 4). Such spectral evolution feature of prompt emission could be explained as photospheric emission is most commonly observed near the trigger time, while synchrotron-like emission becomes more prevalent in later times, indicating the coexistence of multiple emission mechanisms as the burst evolves (Li et al., 2021).
@@ 3.2 Pulse-wise prompt emission correlation and analysis
The pulse-wise analysis of GRB prompt emission provides a valuable perspective on the Amati and Yonetoku correlations (key relationships used to explore the physical properties of GRBs), enhancing our understanding of GRB spectral evolution and energetics on a finer timescale (Chand et al., 2020). By examining each pulse independently, rather than averaging over the entire prompt emission phase, we gain insights into the distinct characteristics of individual emission episodes within a single burst. In the Amati correlation, which relates the isotropic energy () to the spectral peak energy in the rest frame (Amati, 2006), each pulse may fall along or diverge from the established long and short GRB trends, providing information about the pulse-specific energy release mechanisms. Such an analysis not only clarifies the evolutionary trajectory of the GRB emission but also tests the applicability of the Amati correlation at sub-burst levels (Basak & Rao, 2013).
In the top-left panel of Figure 10, we illustrate the positioning of each pulse (plotted as colored squares) of GRB 230204B within the established Amati correlation. For reference, samples of long and short bursts from Minaev & Pozanenko (2020) are represented by orange and blue circles, respectively. The linear fits corresponding to these groups are indicated by solid orange and blue lines, with 3 uncertainties depicted by shaded bands around each fit. The time-integrated and pulse-wise analysis of GRB 230204B follows the Amati plane of long GRB populations. The top-right panel of Figure 10 presents GRB 230204B positions within the Yonetoku correlation framework, a relation between peak energy and the isotropic peak luminosity () in the rest frame (Yonetoku et al., 2004). Here, the short and long GRB samples from Nava et al. (2012) are shown by blue and orange circles, with the 3 scatter around each population’s best-fit line visualized by shaded areas. GRB 230204B follows the Yonetoku correlation and lies on the upper side of this correlation. The results in both correlations offer an insightful examination of GRB 230204B, highlighting the pulse-specific properties within established GRB behavior and indicating the energetic and luminous nature of the burst.
@@ 3.2.1 Spectral Peak-Duration Distribution Analysis
Long-duration GRBs are generally characterized by a softer spectrum relative to the typically harder spectrum observed in the majority of short-duration GRBs, as demonstrated in previous studies (e.g., Kouveliotou et al. 1993). To analyze this characteristic for GRB 230204B, we measured both the time-integrated and episode-wise spectral peak energy (an indicator of spectral hardness) alongside the duration (the time during which 90% of the burst’s total fluence is detected, for example, calculation for the second pulse of GRB 230204B is shown in the bottom left of Figure A1 of the appendix). We then plotted the pulse-wise values on the - plane to compare with a larger dataset of GRBs detected by the Fermi GBM, as illustrated in the bottom right panel of Figure 10.
Our results indicate that the time-integrated and episode-wise properties of GRB 230204B exhibit softer values and durations exceeding 2 seconds, placing it on the right side of the spectral peak-energy versus duration plane. This position aligns well with the properties typically observed for long-duration GRBs, as they occupy the region in the - plane corresponding to longer durations and softer spectra. This comparison confirms that GRB 230204B follows the standard characteristics of long-duration GRBs.

### v2

@@ 2.4.2 BOOTES robotic telescope
BOOTES followed up GRB 230204B with three of the six 60 cm robotic telescopes, namely the ones at the BOOTES-4/MET station in Lijiang Astronomical Observatory (Yunnan, China), at BOOTES-5/JGT station in Observatorio Astronómico Nacional of San Pedro Mártir (Baja California, México) and at the BOOTES-2/TELMA station in La Mayora (Málaga, Spain). Following the MAXI trigger, the BOOTES-4/MET telescope performed two epoch observations at 2023-02-04 22:19:12 UT and 2023-02-05 17:42:04 UT. The BOOTES-2/TELMA telescope and the BOOTES-5/JGT telescope also executed follow-up observations at 2023-02-05 04:16:55 UT and 2023-02-05 07:43:45 UT, respectively. A series of images were obtained using the clear filter with exposures of 1 sec, 5 sec, and 10 sec at BOOTES-4/MET telescope during the first epoch, and the optical afterglow was clearly detected during the decaying phase. However, during the subsequent multi-filter monitoring (Sloan- and clear filters, 60 sec exposure), the optical afterglow was not detected on the stacked images for each epoch. This non-detection is likely due to the fact that the late-time stacked images were not deep enough to detect the rapidly fading afterglow. All the images were processed with bias, dark subtraction, and flat-field correction using custom IRAF routines (2021MNRAS.505.4086G). Photometry was carried out using the standard IRAF package, and the images were calibrated with nearby comparison stars from the USNO-B1.0 (for clear filter) and SDSS (for filters) catalogs.
@@ 2.4.3 DOT
We initiated observations using the DOT (2016RMxAC..48...83P; 2024BSRSL..93..683G) on 2023-02-06 at 01:49:47.939 UT, approximately 1.16 days after the trigger. Multiple frames with an exposure time of 180 sec each were observed in the , , and bands, and continuous observations were conducted over four consecutive nights using the same filters. For image preprocessing, we employed the standard IRAF framework, and DAOPHOT II was used to determine magnitudes from the cleaned files. We stacked the images for each night in the same filter to increase the signal-to-noise ratio. PSF photometry was performed on the final stacked images (2022JApA...43...82G; 2022JApA...43...11G). Due to the bright moon phase during the burst observation, we only detected GRB 230204B in the and -band images from the first night. For the remaining filters, we could only establish upper limits. Photometric calibration is performed using the standard stars from the Pan-STARRS catalog.
@@ 2.4.4 Swift/UVOT
The Swift/UVOT began settled observations of the field of GRB 230204B 80.6 ks after the MAXI trigger (2023GCN.33265....1S). No optical afterglow was detected by UVOT. Photometry was obtained from co-added UVOT image mode data. The source counts were extracted from the UVOT stacked images using a source region of 5 arcsec radius. Background counts were extracted using an aperture of 20 arcsec in radius located in a source-free region near the GRB. We use the Swift tool uvotsource to obtain the background-subtracted count rates using source and background apertures of 5 and 20 arcseconds in radius, respectively. They were converted to magnitude using the UVOT photometric zero points (bre11; poole).
@@ 2.4.5 Radio Emission
We observed GRB 230204B on 8 February 2023 (3.6 days post-burst), with the Australia Telescope Compact Array (ATCA, Project code CX527) at a wide range of frequencies centered on 5.5, 9.0, 16.7, 21.2, 33, and 35 GHz, each with a 2048 MHz-wide band. We reduced the visibility data using standard routines in MIRIAD (1995ASPC...77..433S). We used a combination of manual and automatic RFI flagging before calibration, conducted with MIRIAD tasks uvflag and pgflag, respectively. We used PKS 193463 to determine the bandpass response for frequency bands 5.5 and 9 GHz, and PKS 0727115 as the bandpass calibrator for 16.7, 21.2, 33, and 35 GHz bands. We used PKS 193463 to calibrate the flux density scale for all frequency bands. We used PKS 1256220 to calibrate the time-variable complex gains for all epochs and frequency bands. After calibration, we inverted the visibilities using a robust weighting of 0.5 and then used the CLEAN algorithm (1980A&A....89..377C) on the target source field using standard MIRIAD tasks INVERT, CLEAN, and RESTOR to obtain the final images.
We detect the radio counterpart at 16.7 and 21.2 GHz at a position consistent with the BOOTES/MASTER/GIT optical counterpart position (2023GCN.33269....1S). For all other frequencies (5.5, 9.0, 33.0, 35.0, 0.842, and 0.943 GHz), we report 5-sigma upper limits (see Table 8 of the appendix). The GRB position was also observed with the Australian SKA Pathfinder by the Rapid ASKAP Continuum Survey (RACS; 2020PASA...37...48M) at 0.842, and 0.943 GHz on 2023 July 1 and 2024 Jan 5. We find no radio detections in preliminary data and report 5-sigma upper limits of 1.5 mJy at 0.842 and 0.943 GHz. There were also no archival radio detection pre-burst within 1 arcminute in surveys: the National Radio Astronomy Observatory VLA Sky Survey (1998AJ....115.1693C) at 1.4 GHz, the Sydney University Molonglo Sky Survey (2003MNRAS.342.1117M) at 843 MHz, the Very Large Array Sky Survey (2020PASP..132c5001L) at 3 GHz or the RACS survey. We measured a VLASS 5-sigma upper limit of 0.69 mJy at 3 GHz.
@@ 2.4.6 Swift BAT Survey Data
Swift/XRT did not observe the location of X-ray afterglow of GRB 230204B until 91.5 ks post-burst. Therefore, to search for the hard X-ray early afterglow, we used the BatAnalysis python package (BatAnalysis) to analyze Swift BAT survey data from 2023-02-03 to 2023-02-11. We analyzed survey data, where the location of the GRB had at least a partial coding fraction of on the BAT detector plane. As expected, the GRB afterglow was not detected in the BAT survey data, which is typical since hard X-ray afterglows have never been detected at such late times. However, we were able to place 5 upper limits on the emission in the 14-195 keV energy band. In Figure 3, we show the flux upper limits derived from individual BAT survey snapshots, which correspond to shorter time bins and thus shallower limits. To obtain more stringent constraints, we also performed a mosaiced analysis using 1-day bins for the two days after the trigger, yielding deeper upper limits of erg sec-1 cm-2. Although the BAT survey provides the deepest available constraints in the 14–195 keV range during this period, the obtained limits are not sufficiently deep to impose meaningful physical constraints on the afterglow. In particular, the derived upper limits are significantly above the flux level expected from a simple extrapolation of the contemporaneous Swift/XRT spectrum to hard X-rays. Thus, while these upper limits confirm the absence of an unusually bright hard X-ray component, they are too shallow to provide additional constraints on the standard afterglow emission.
@@ 3 Results
3.1 Prompt emission temporal and spectral characterization
For the analysis of Fermi GBM data, we utilized the Fermi GBM Data Tools222https://fermi.gsfc.nasa.gov/ssc/data/analysis/gbm/gbm_data_tools/gdt-docs/ to extract temporal information. We selected specific NaI (NaI 7 and NaI 8) and BGO (BGO 1) detectors for the analysis based on their optimal viewing angles relative to the GRB’s position to maximize the signal-to-noise ratio. For the NaI detectors, we focused on two energy bands: 8-50 keV and 50-300 keV, allowing us to explore both low and high-energy photon populations. The 8-50 keV band was chosen to provide a more comprehensive coverage of the soft X-ray emission, while the 50-300 keV band captures the higher-energy component. For BGO 1, two bands were used: 0.3-1 MeV and 1-40 MeV, to investigate higher-energy photons. The prompt emission background-subtracted light curve of GRB 230204B, obtained using Fermi GBM across different energy channels, reveals a complex multi-pulse structure, as shown in Figure 4. This GRB exhibits four distinct emission episodes, each separated by periods of quiescence, with the shaded regions highlighting the intervals selected for time-integrated spectral analysis. The time windows (pulse 1: T0- 0.80 to T0 + 4.97 sec, pulse 2: T0 + 37.00 to T0 + 68.15 sec, pulse 3: T0 + 109.45 to T0 + 161.90 sec, and pulse 4: T0 + 172.58 to T0 + 229.69 sec) for these intervals were determined using the Bayesian Block algorithm, which adaptively identifies significant temporal structures in the light curve based on statistical criteria. The energy-dependent light curves show that the temporal structure and intensity vary significantly across different energy bands, reinforcing the diverse nature of GRB light curves. Notably, the hardness ratio (HR) demonstrates evolution over time, indicating spectral changes during the burst. The inset plot shows the MAXI light curve, covering the 2-20 keV energy range, further subdivided into 2-4 keV, 4-10 keV, and 10-20 keV. Although MAXI detected only part of the GRB, its observations provide valuable insights into the lower-energy emission and suggest a much softer nature of the last pulse of GRB 230204B.
We analyzed the time-integrated (T0-0.80 to T0 +229.69 sec) Fermi spectrum of GRB 230204B using the Multi-Mission Maximum Likelihood framework (3ML) to evaluate its spectral characteristics. We explore several phenomenological models, including the Band function, Cutoff power-law (CPL), and combinations with a Blackbody component, which is theoretically expected to originate from the photospheric emission of GRBs (2010ApJ...709L.172R; 2013ApJ...765..103L). For each model, we employed Bayesian spectral fitting to obtain the Deviance Information Criterion (DIC) values, allowing us to systematically compare the model fits. The results indicate that the Band function provides a statistically better fit to the time-integrated spectrum compared to the CPL alone, as indicated by the DIC values. Furthermore, the addition of a Blackbody component alongside the Band function yielded an even lower DIC value, suggesting that this model combination better describes the observed spectral features.
@@ 3.1 Prompt emission temporal and spectral characterization
GRBs are known for their significant spectral evolution and high variability, with emission properties often varying notably across individual pulses within multi-pulse bursts. Liang2021 conduct a comprehensive time-resolved Bayesian spectral analysis of well-defined GRB pulses observed by Fermi. They observe a trend toward spectral softening over time, largely reflected in the gradual decrease of the low-energy power-law index. We compared the time-resolved spectral analysis of GRB 230204B with the sample studied by Liang2021. We also observe that the low-energy power-law indices of GRB 230204B become softer over time (hard to soft), with a distinct spectrum evolution from the first to later episodes (see Figure 5). Such a spectral evolution feature of prompt emission could be explained as photospheric emission is most commonly observed near the trigger time, while synchrotron-like emission becomes more prevalent in later times, indicating the coexistence of multiple emission mechanisms as the burst evolves (Liang2021).
@@ 3.2 Pulse-wise prompt emission correlation and analysis
The pulse-wise analysis of GRB prompt emission provides a valuable perspective on the Amati and Yonetoku correlations (key relationships used to explore the physical properties of GRBs), enhancing our understanding of GRB spectral evolution and energetics on a finer timescale (2020ApJ...898...42C). By examining each pulse independently, rather than averaging over the entire prompt emission phase, we gain insights into the distinct characteristics of individual emission episodes within a single burst. In the Amati correlation, which relates the isotropic energy () to the spectral peak energy in the rest frame (2006MNRAS.372..233A), each pulse may fall along or diverge from the established long and short GRB trends, providing information about the pulse-specific energy release mechanisms. Such an analysis not only clarifies the evolutionary trajectory of the GRB emission but also tests the applicability of the Amati correlation at sub-burst levels (2013MNRAS.436.3082B).
In the top-left panel of Figure 11, we illustrate the positioning of each pulse (plotted as colored squares) of GRB 230204B within the established Amati correlation. For reference, samples of long and short bursts from 2020MNRAS.492.1919M are represented by orange and blue circles, respectively. The linear fits corresponding to these groups are indicated by solid orange and blue lines, with 3 uncertainties depicted by shaded bands around each fit. The time-integrated ( = 1.92 erg) shows that GRB 230204B follows the Amati plane of long GRB populations. However, when considering the episodes individually, we find that while later pulse Episode 4 remains broadly consistent with the long GRB population, Episodes 1, 2, and 3 fall into an intermediate region between the long and short GRB distributions, indicating a less well-defined classification.
The top-right panel of Figure 11 presents GRB 230204B positions within the Yonetoku correlation framework, a relation between peak energy and the isotropic peak luminosity () in the rest frame (2004ApJ...609..935Y). Here, the short and long GRB samples from 2012MNRAS.421.1256N are shown by blue and orange circles, with the 3 scatter around the population’s best-fit line visualized by a shaded area. Similar to the Amati plane, the overall burst lies on the upper side of the Yonetoku correlation for long GRBs.
@@ 3.2.1 Spectral Peak-Duration Distribution Analysis
Long-duration GRBs are generally characterized by a softer spectrum relative to the typically harder spectrum observed in the majority of short-duration GRBs, as demonstrated in previous studies (e.g., 1993ApJ...413L.101K). To analyze this characteristic for GRB 230204B, we measured both the time-integrated and episode-wise spectral peak energy (an indicator of spectral hardness) alongside the duration (the time during which 90% of the burst’s total fluence is detected, values are 214.02 sec, 4.86 sec, 24.32 sec, 41.98 sec, and 51.46 sec for time-integrated duration, episode 1, episode 2, episode 3, and episode 4, respectively). We then plotted the pulse-wise values on the - plane to compare with a larger dataset of GRBs detected by the Fermi GBM, as illustrated in the bottom panel of Figure 11.
Our results show that the time-integrated properties of GRB 230204B exhibit a relatively soft value and duration longer than 2 sec, placing the burst clearly within the Type II (collapsar) GRB population. However, the episode-wise analysis reveals interesting diversity in spectral and temporal properties: Episode 1 lies an intermediate region in both the - and Amati correlation (see Section 3.2) plane, with relatively short duration ( 4.86 sec), while Episodes 2 and 3 show progressive evolution toward longer durations and softer spectra. Episode 4 firmly places itself within the long GRB distribution with the softest spectrum ( -0.80) and longest duration (T90 51.46 sec). This diversity likely reflects evolving emission mechanisms throughout the burst. The spectral evolution from photosphere-dominated to synchrotron-dominated emission supports hybrid jet models where thermal and non-thermal components coexist (Liang2021). Alternatively, for collapsar with residual stellar envelopes, the jet’s interaction with material at different radii can produce distinct emission episodes with varying durations and spectral properties (2003ApJ...584..390W; 2003ApJ...586..356Z).
Our analysis demonstrates that individual emission episodes within a single long GRB can exhibit durations and spectral properties overlapping with the short GRB population and provides evidence for complex, evolving emission physics throughout the burst duration. This finding adds to growing evidence that the traditional 2-second duration threshold for GRB classification may be overly simplistic and that progenitor identification requires consideration of additional properties beyond alone (2025arXiv250800142G). GRB 230204B exemplifies how such multi-layered approaches reveal complex emission physics that would remain hidden in time-integrated analyses alone. This underscores the importance of pulse-resolved studies for understanding the diversity of GRB emission mechanisms and jet physics.